A survey shows that 10% of students are victimized by bullies in school each year. A random sample of 527 students shows a victimization rate of 14%. Are students more likely to be bullied in school? Assume significance level at 0.01.


1
Expert's answer
2021-12-23T12:28:42-0500

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\leq 0.1"

"H_1: p>0.1"

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z > 2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p}{\sqrt{\dfrac{p(1-p)}{n}}}=\dfrac{0.14-0.1}{\sqrt{\dfrac{0.1(1-0.1)}{527}}}"

"\approx3.060864"

Since it is observed that "z = 3.060864 >2.3263= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>3.060864)\approx0.0011035," and since "p=0.0011035<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than "0.1," at the "\alpha = 0.01" significance level.